# Structure-preserving contrastive learning for spatial time series

### v2

@@ Abstract
Informative representations enhance model performance and generalisability in downstream tasks. However, learning self-supervised representations for spatially characterised time series, like traffic interactions, poses challenges as it requires maintaining fine-grained similarity relations in the latent space. In this study, we incorporate two structure-preserving regularisers for the contrastive learning of spatial time series: one regulariser preserves the topology of similarities between instances, and the other preserves the graph geometry of similarities across spatial and temporal dimensions. To balance contrastive learning and structure preservation, we propose a dynamic mechanism that adaptively weighs the trade-off and stabilises training. We conduct experiments on multivariate time series classification, as well as macroscopic and microscopic traffic prediction. For all three tasks, our approach preserves the structures of similarity relations more effectively and improves state-of-the-art task performances. The proposed approach can be applied to an arbitrary encoder and is particularly beneficial for time series with spatial or geographical features. Furthermore, this study suggests that higher similarity structure preservation indicates more informative and useful representations. This may help to understand the contribution of representation learning in pattern recognition with neural networks. Our code is made openly accessible with all resulting data at https://github.com/yiru-jiao/spclt.
@@
Contrastive learning , similarity preservation , time series , spatio-temporal data , traffic interaction
@@ 1 Introduction
Self-supervised representation learning (SSRL) theoretically can learn latent embeddings that facilitate downstream tasks [1, 2, 3, 4]. Also, it is practically shown to improve model generalisability [5, 6]. The latter is in particular valuable for real-world applications, where both measurements and labels are often uncertain and unreliable. In fact, SSRL has been widely applied across fields such as computer vision, natural language processing, and recommendation systems. There are many literature reviews, to name a few, [7, 8, 9].
@@
In SSRL of time series, contrastive learning is becoming the mainstay technique, contributing to applications in, e.g., anomaly detection [10, 11], robotics [12, 13], and healthcare [14]. This widespread adoption is supported by empirical investigation. In 2022, Lafabregue et al. [15] conducted an extensive experimental comparison over 300 combinations of network architectures and loss functions to evaluate the performance of time series representation learning. One of their key findings is that the reconstruction loss used by traditional autoencoders does not sufficiently fit temporal patterns. Instead, contrastive learning has emerged as a more effective approach, which embeds similar samples closer together while dissimilar samples farther apart in the latent space [16, 17].
@@
Unique challenges arise when learning contrastive representations for spatial time series. First, data with both temporal and spatial characteristics demand more fine-grained similarity comparisons, which underpins contrastive learning. Financial time series may be considered similar even if some variables show significant divergence, while movement traces with very different spatial features can be anything but similar. Second, effective representation of spatial time series needs to capture spatio-temporal patterns at the certain scale required by a practical task. For example, traffic interactions involve two different spatial scales: at the macroscopic scale, traffic flow measures collective road usage evolving over the road network; at the microscopic scale, trajectories describe the motion dynamics of individual road users such as car drivers, cyclists, and pedestrians, in local road space.
@@
To address the challenges, we incorporate two regularisers at different scales to preserve the original similarity structure in the latent space for time series contrastive learning. One is a topology-preserving regulariser for the global scale, and the other is a graph-geometry-preserving regulariser for the local scale. This incorporation can be simplified as a weighted loss ${\mathcal{L}}=\eta_{\text{CLT}}\cdot\ell_{\text{CLT}}+\eta_{\text{SP}}\cdot\ell_{\text{SP}}+r_{\bm{\eta}}$, where we propose a mechanism to dynamically balance the weights $\eta_{\text{CLT}}$ of contrastive learning for time series (CLT) and $\eta_{\text{SP}}$ of structure preservation (SP). Within this mechanism, the adaptive trade-off between contrastive learning and structure preservation is based on the uncertainties of their corresponding terms $\ell_{\text{CLT}}$ and $\ell_{\text{SP}}$; meanwhile, the term $r_{\bm{\eta}}$ adds regularisation against overfitting of the dynamic weights.
@@
The proposed approach is applicable to spatial time series in general, and we consider traffic interaction as a specific case in this paper. To validate the approach, we conduct experiments on tasks of 1) multivariate time series classification, where we benchmark against the current state-of-the-art (SOTA) models, i.e., [18] and [19]; and 2) traffic prediction, where we use [20] for macroscopic benchmark and [21] for microscopic. Along these experiments, the efficiency of this approach is evaluated with various model architectures. In addition to performance improvement, we also investigate and discuss the importance of preserving similarity structure. Below is a summary of the key contributions in this study.
@@
  • We introduce an approach that incorporates structure-preserving regularisation in contrastive learning of multivariate time series, to maintain finer-grained similarity relations in the latent space of sample representations. We propose a dynamic weighing mechanism to adaptively balance contrastive learning and structure preservation during training. This approach can be applied to an arbitrary encoder for more effective self-supervised representation learning.
@@
  • Considering neural network pattern recognition as learning the conditional probability distribution of outputs over inputs, the similarity structure hidden in the input data implies the distribution of conditions. Our approach is therefore important to preserve the original distribution in the latent space for more effective representational learning. This is particularly beneficial for time-series data with spatial or geographical features as they require finer-grained information, such as in robotics, meteorology, remote sensing, urban planning, etc.
The rest of this paper is organised as follows. In Section 2, we briefly review related works in the literature. We use Section 3 to systematically introduce the methods. Then in Section 4, we present the demonstration experiments and according results. With Section 5, we discuss the importance of preserving similarity structure for representation learning. Finally, Section 6 concludes this study.
@@ 4.2 Multivariate time series classification
Table 4 displays the classification performance on spatial and non-spatial UEA datasets. Next to the averaged accuracy, we also include the loss values on test sets to offer more information. More detailed results can be found in Tables B1 and B2 in B, where we present the classification accuracy with different representation learning losses for each dataset. Then we use Table 4 to more specifically compare the relative improvements induced by adding a topology or graph-geometry preserving regulariser. The improvement is measured by the percentage of accuracy difference from the corresponding baseline performance.
@@
Tables 4 and 4 clearly show that structure preservation improves classification accuracy, not only when time series data involve spatial features, but also when they do not. The relative improvements in Table 4 are higher for non-spatial datasets than for spatial datasets, which is because the datasets without spatial features are more difficult to learn in the UEA archive. As is shown in Table 4, the loss of contrastive learning decreases when a structure-preserving regulariser is added for spatial datasets, while increases for non-spatial datasets. This implies that preserving similarity structure is well aligned with contrastive learning for spatial datasts, and can even enhance contrastive learning.
@@ 4.3 Macroscopic and microscopic traffic prediction
Table 7 then displays the corresponding evaluation on similarity structure preservation, which is obtained by assessing the encoders after fine-tuning for traffic prediction. The results show that the better performing methods in macro-traffic prediction (i.e., GGeo-SoftCLT, Topo-SoftCLT) and micro-traffic prediction (i.e., Topo-TS2Vec, GGeo-TS2Vec) preserve more similarity structures at both global and local scales. In general, the metric values are close for the same task across methods; however, when a method has a significant advantage over the others, it indicates superior performance. Examples for macroscopic traffic prediction are Topo-SoftCLT in local Cont. and GGeo-SoftCLT in global dRMSE; for microscopic traffic prediction are Topo-TS2Vec in local Trust. and GGeo-TS2Vec in global Cont.
@@ 4.4 Training efficiency
a Detailed results are referred to Tables B3 and B4 in B.
@@
Figure 1: Fine-tuning progress of models trained from scratch and pretrained with different losses in macroscopic traffic prediction. Values of the final performance are referred to Table 6 for DGCN, to Tables C1 and C2 in C for LSTM and GRU.
For all models of DGCN, LSTM, and GRU, structure preservation consistently enhances prediction performance compared to training from scratch (No pretraining). The enhancement is significant when using LSTM and GRU, achieving 6.68% and 10.14% improvement in RMSE, respectively. Meanwhile, the progress of structure preservation is stable when the encoder is pretrained, and maintains the advantage over no pretraining throughout the fine-tuning process. In contrast, for DGCN which is a more sophisticated model tailored for the task, training from scratch is already very effective and pretraining brings relatively minor improvement. This implies that certain model architectures are better suited to a specific task than others, and the preservation of similarity structure in the latent representations may be a good indicator for model selection.
@@ 5 Discussion
A key consideration when applying our approach lies in selecting an appropriate loss function. This involves two layers of choices: TS2Vec versus SoftCLT, and topology-preserving (Topo-) versus graph-geometry-preserving (GGeo-) regularisation. For the first choice, TS2Vec is generally more suitable for classification tasks with fewer classes as TS2Vec only compares the similarity between two different augmentations of the same sample. In contrast, SoftCLT incorporates all samples in a mini-batch by assigning soft labels based on similarity. This performs a more detailed similarity comparison and thus is advantageous for tasks with a larger number of classes or for regression. As shown in Figure 2, in the UEA archive, datasets for which (Topo/GGeo-)TS2Vec achieves the best accuracy tend to have fewer classes than those with the best performance achieved by (Topo/GGeo-)SoftCLT. In essence, SoftCLT implicitly embeds the similarity structure through soft labels.
@@ 6 Conclusion
This paper presents an approach to structure-preserving contrastive learning for spatial time series, where we propose a dynamic mechanism to adaptively balance contrastive learning and structure preservation. Extensive experiments demonstrate that our methods can improve the SOTA performance, including for multivariate time series classification in various contexts and for traffic prediction at both macroscopic and microscopic scales. In particular, preserving similarity structure is shown to be well aligned with and beneficial to contrastive learning for spatio-temporal data. This provides a solid yes to the proposed question: preserving fine-grained structure of similarity relations in contrastive learning can enhance downstream task performance for spatial time series.
Our findings emphasise that preserving the original similarity structure in the latent space is crucial for effective representational learning, as it impacts a neural network’s ability to model the underlying conditional distribution. In general, adding structure-preserving regularisation has a limited impact on the efficiency of self-supervised representation learning. It can be computationally intensive when the time sequence is long; however, evident performance improvement makes it an acceptable price to pay for utilising the information embedded in time series data. In addition, our experiments suggest that higher similarity structure preservation may be a good indicator of more informative representations, highlighting that the structural information of similarities in spatio-temporal data remains yet to be exploited. Given that many real-world practices involve spatial time series, this study and future research based on it can be applied not only to traffic interactions, but also to any that can benefit from preserving specific structures in similarity relations.
@@ Appendix B Detailed results on UEA datasets
This section provides detailed comparisons of evaluation results for the used 28 datasets in the UEA archive. Tables B1 and B2 present the results of classification accuracy. Tables B3 and B4 present the training time for self-supervised representation learning.
Table B1: Detailed evaluation of classification accuracy on spatial datasets in the UEA archive.
@@
Table B2: Detailed evaluation of classification accuracy on non-spatial datasets in the UEA archive.
@@
Table B3: Detailed representation training time per epoch (unit: s) on spatial datasets in the UEA archive.
@@
Table B4: Detailed representation training time per epoch (unit: s) on non-spatial datasets in the UEA archive.
@@
In addition, to visually show the effect of differently regularised contrastive learning losses on representation, we apply t-SNE to compress the encoded representations into 3 dimensions, as plotted in Figure B2 for the dataset Epilepsy, and Figure B2 for RacketSports. The classes are indicated by colours. We use these two datasets because they are visualisation-friendly, with 4 classes and around 150 test samples.
Figure B1: Encoded representations after training with different losses on the test set of Epilepsy.
Figure B2: Encoded representations after training with different losses on the test set of RacketSports.
@@ Appendix C Detailed results of macroscopic prediction with LSTM and GRU
This section provides additional tables presenting the evaluation of the final results using LSTM and GRU in macroscopic traffic prediction. Table C1 shows the task-specific metrics and Table C2 shows the metrics for global structure preservation.
Table C1: Macroscopic traffic prediction evaluation with LSTM and GRU encoders.
@@
Table C2: Global structure preservation of LSTM and GRU encoders in macroscopic traffic prediction task.

### v3

@@ Abstract
Neural network models are increasingly applied in transportation research to tasks such as prediction. The effectiveness of these models largely relies on learning meaningful latent patterns from data, where self-supervised learning of informative representations can enhance model performance and generalisability. However, self-supervised representation learning for spatially characterised time series, which are ubiquitous in transportation domain, poses unique challenges due to the necessity of maintaining fine-grained spatio-temporal similarities in the latent space. In this study, we introduce two structure-preserving regularisers for the contrastive learning of spatial time series: one regulariser preserves the topology of similarities between instances, and the other preserves the graph geometry of similarities across spatial and temporal dimensions. To balance the contrastive learning objective and the need for structure preservation, we propose a dynamic weighting mechanism that adaptively manages this trade-off and stabilises training. We validate the proposed method through extensive experiments, including multivariate time series classification to demonstrate its general applicability, as well as macroscopic and microscopic traffic prediction to highlight its particular usefulness in encoding traffic interactions. Across all tasks, our method preserves the similarity structures more effectively and improves state-of-the-art task performances. This method can be integrated with an arbitrary neural network model and is particularly beneficial for time series data with spatial or geographical features. Furthermore, our findings suggest that well-preserved similarity structures in the latent space indicate more informative and useful representations. This provides insights to design more effective neural networks for data-driven transportation research. Our code is made openly accessible with all resulting data at https://github.com/yiru-jiao/spclt.
@@
Contrastive learning , representation learning , time series , spatio-temporal data , traffic interaction
@@ 1 Introduction
Modern transportation systems create massive streams of spatially distributed time series data, such as traffic speeds across road networks and transit ridership through various locations over time. Extracting useful patterns from these data is crucial, especially as neural network models are increasingly used for a range of downstream tasks such as traffic forecasting [1, 2], congestion detection [3, 4], and mobility analysis [5, 6]. In recent years, self-supervised representation learning (SSRL) has emerged as a promising approach to leverage such large-scale datasets [7]. By learning informative latent representations, SSRL can effectively facilitate model performance [8, 9, 10] and generalisability [11, 12] in downstream tasks. This advantage is especially valuable in transportation research, where real-world sensor measurements and labels are often noisy or sparse.
@@
In SSRL of time series, contrastive learning is becoming the mainstay technique. This adoption is supported by empirical investigation. In 2022, Lafabregue et al. [13] conducted an extensive experimental comparison over 300 combinations of network architectures and loss functions to evaluate the performance of time series representation learning. One of their key findings is that the reconstruction loss used by traditional autoencoders does not sufficiently fit temporal patterns. Instead, contrastive learning has emerged as a more effective approach, which explicitly pulls similar instances closer and pushes dissimilar instances farther apart in the latent space of representations [14, 15]. This mechanism encourages neural networks to organise the latent space according to the inherent similarities in data, yielding representations that capture meaningful patterns.
@@
Unique challenges arise when learning contrastive representations for spatially characterised time series data. A foremost difficulty is the need to preserve fine-grained similarity structures among data instances in the latent space. The notion of similarity for spatial time series can be subtle and highly domain-specific. For example, financial time series may be considered similar even if some variables show significant divergence, while movement traces with very different spatial features can be anything but similar. Beyond preserving fine-grained similarities, spatially characterised time series such as traffic interactions can involve multiple scales of spatio-temporal patterns. At the macroscopic scale, traffic flow measures collective road usage evolving over the road network; at the microscopic scale, trajectories describe the motion dynamics of individual road users such as car drivers, cyclists, and pedestrians, in local road space. SSRL for spatial time series must accommodate such heterogeneity, capturing patterns at the appropriate level of granularity for the targeted task.
@@
To address these challenges, this study explores contrastive learning regularised by structure preservation to better capture the subtle similarities in spatial time series data. We introduce two regularisers at different scales to preserve the original similarity structure in the latent space. One is a topology-preserving regulariser for the global scale, and the other is a graph-geometry-preserving regulariser for the local scale. This incorporation can be simplified as a weighted loss ${\mathcal{L}}=\eta_{\text{CLT}}\cdot\ell_{\text{CLT}}+\eta_{\text{SP}}\cdot\ell_{\text{SP}}+r_{\bm{\eta}}$, where we propose a mechanism to dynamically balance the weights $\eta_{\text{CLT}}$ of contrastive learning for time series (CLT) and $\eta_{\text{SP}}$ of structure preservation (SP). Within this mechanism, the adaptive trade-off between contrastive learning and structure preservation is based on the uncertainties of their corresponding terms $\ell_{\text{CLT}}$ and $\ell_{\text{SP}}$; meanwhile, the term $r_{\bm{\eta}}$ adds regularisation against overfitting of the dynamic weights.
@@
The proposed method is applicable to spatial time series in general, while we highlight its particular usage for traffic interactions in this paper. To thoroughly validate the method, we conduct experiments on tasks of 1) multivariate time series classification, where we benchmark against the current state-of-the-art (SOTA) models, i.e., [16] and [17]; and 2) traffic prediction, where we use [18] for macroscopic benchmark and [19] for microscopic. Along with these experiments, the efficiency of this method is evaluated with various network architectures. In addition to performance improvement, we also investigate and discuss the importance of preserving similarity structure. Below is a summary of the key contributions in this study.
@@
  • We introduce a method that incorporates structure-preserving regularisation in contrastive learning of multivariate time series, to maintain finer-grained similarity structures in the latent space of sample representations. We propose a dynamic weighting mechanism to adaptively balance contrastive learning and structure preservation during training. This method can be applied to an arbitrary neural network model for more effective representation learning.
@@
  • Considering neural network modelling as learning the conditional probability distribution of outputs over inputs, the similarity structure hidden in the input data implies the distribution of conditions. Our method is therefore important to preserve the original distribution in the latent space for more effective representational learning. This is particularly beneficial for time series data with spatial or geographical features as they require finer-grained information, not only in the field of transportation, but also robotics, meteorology, remote sensing, etc.
The rest of this paper is organised as follows. In Section 2, we briefly review related work in the literature. We use Section 3 to systematically introduce the methods. Then we explain the demonstration experiments in Section 4 and present according results in Section 5. With Section 6, we discuss the importance of preserving similarity structure for representation learning. Finally, Section 7 concludes this study.
@@ 5.1 Multivariate time series classification
Table 4 displays the classification performance on spatial and non-spatial UEA datasets. Next to the averaged accuracy, we also include the loss values on test sets to offer more information. More detailed results can be found in Tables B.1 and B.2 in B, where we present the classification accuracy with different representation learning losses for each dataset. Then we use Table 4 to more specifically compare the relative improvements induced by adding a topology or graph-geometry preserving regulariser. The improvement is measured by the percentage of accuracy difference from the corresponding baseline performance.
@@
Tables 4 and 4 clearly show that structure preservation improves classification accuracy, not only when time series data involve spatial features, but also when they do not. The relative improvements in Table 4 are higher for non-spatial datasets than for spatial datasets, which is because the datasets without spatial features are more difficult to learn in the UEA archive. As is shown in Table 4, the loss of contrastive learning decreases when a structure-preserving regulariser is added for spatial datasets, while increases for non-spatial datasets. This implies that preserving similarity structure is well aligned with contrastive learning for spatial datasets, and can even enhance contrastive learning.
@@ 5.2 Macroscopic and microscopic traffic prediction
Table 7 then displays the corresponding evaluation on similarity structure preservation, which is obtained by assessing the encoders after fine-tuning for traffic prediction. The results show that the better-performing methods in macro-traffic prediction (i.e., GGeo-SoftCLT, Topo-SoftCLT) and micro-traffic prediction (i.e., Topo-TS2Vec, GGeo-TS2Vec) preserve more similarity structures at both global and local scales. In general, the metric values are close for the same task across methods; however, when a method has a significant advantage over the others, it indicates superior performance. Examples for macroscopic traffic prediction are Topo-SoftCLT in local Cont. and GGeo-SoftCLT in global dRMSE; for microscopic traffic prediction are Topo-TS2Vec in local Trust. and GGeo-TS2Vec in global Cont.
@@ 5.3 Training efficiency
a Detailed results are referred to Tables B.3 and B.4 in B.
@@
Figure 1: Fine-tuning progress of models trained from scratch and pretrained with different losses in macroscopic traffic prediction. Values of the final performance are referred to Table 6 for DGCN, to Tables C.1 and C.2 in C for LSTM and GRU.
For all models of DGCN, LSTM, and GRU, structure preservation consistently enhances prediction performance compared to training from scratch (No pretraining). The enhancement is significant when using LSTM and GRU, achieving 6.68% and 10.14% improvement in RMSE, respectively. Meanwhile, the progress of structure preservation is stable when the encoder is pretrained, and maintains the advantage over no pretraining throughout the fine-tuning process. In contrast, for DGCN, which is a more sophisticated model tailored for the task, training from scratch is already very effective and pretraining brings relatively minor improvement. This implies that certain model architectures are better suited to a specific task than others, and the preservation of similarity structure in the latent representations may be a good indicator for model selection.
@@ 6 Discussion
A key consideration when applying our method lies in selecting an appropriate loss function. This involves two layers of choices: TS2Vec versus SoftCLT, and topology-preserving (Topo-) versus graph-geometry-preserving (GGeo-) regularisation. For the first choice, TS2Vec is generally more suitable for classification tasks with fewer classes as TS2Vec only compares the similarity between two different augmentations of the same sample. In contrast, SoftCLT incorporates all samples in a mini-batch by assigning soft labels based on similarity. This performs a more detailed similarity comparison and thus is advantageous for tasks with a larger number of classes or for regression. As shown in Figure 2, in the UEA archive, datasets for which (Topo/GGeo-)TS2Vec achieves the best accuracy tend to have fewer classes than those with the best performance achieved by (Topo/GGeo-)SoftCLT. In essence, SoftCLT implicitly embeds the similarity structure through soft labels.
@@ 7 Conclusion
This paper presents a method for structure-preserving contrastive learning for spatial time series, where we propose a dynamic mechanism to adaptively balance contrastive learning and structure preservation. This method is generally applicable to time series data with spatial or geographical features, which are particularly abundant in transportation systems. Extensive experiments demonstrate that our methods can improve the SOTA performance, including for multivariate time series classification in various contexts and for traffic prediction at both macroscopic and microscopic scales. Notably, preserving the original similarity structure in the latent space is shown to be well aligned with and beneficial to contrastive learning for spatio-temporal data. This is crucial for effective representational learning from large-scale data, as it impacts a neural network’s ability to model the underlying conditional distribution.
In general, adding structure-preserving regularisation has a limited impact on the efficiency of self-supervised representation learning, while facilitating the convergence of downstream tasks. It can be computationally intensive when the time sequence is long; however, evident performance improvement makes it an acceptable price to pay for utilising the information embedded in time series data. In addition, our experiments suggest that higher similarity structure preservation may be a good indicator of more informative representations, highlighting that the structural information of similarities in spatio-temporal data remains yet to be exploited. Given that many real-world practices involve spatial time series, this study and future research based on it can be applied not only to traffic interactions, but also to any that can benefit from preserving specific structures in similarity relations.
@@ Appendix B Detailed results on UEA datasets
This section provides detailed comparisons of evaluation results for the used 28 datasets in the UEA archive. Tables B.1 and B.2 present the results of classification accuracy. Tables B.3 and B.4 present the training time for self-supervised representation learning.
Table B.1: Detailed evaluation of classification accuracy on spatial datasets in the UEA archive.
@@
Table B.2: Detailed evaluation of classification accuracy on non-spatial datasets in the UEA archive.
@@
Table B.3: Detailed representation training time per epoch (unit: s) on spatial datasets in the UEA archive.
@@
Table B.4: Detailed representation training time per epoch (unit: s) on non-spatial datasets in the UEA archive.
@@
In addition, to visually show the effect of differently regularised contrastive learning losses on representation, we apply t-SNE to compress the encoded representations into 3 dimensions, as plotted in Figure B.2 for the dataset Epilepsy, and Figure B.2 for RacketSports. The classes are indicated by colours. We use these two datasets because they are visualisation-friendly, with 4 classes and around 150 test samples.
Figure B.1: Encoded representations after training with different losses on the test set of Epilepsy.
Figure B.2: Encoded representations after training with different losses on the test set of RacketSports.
@@ Appendix C Detailed results of macroscopic prediction with LSTM and GRU
This section provides additional tables presenting the evaluation of the final results using LSTM and GRU in macroscopic traffic prediction. Table C.1 shows the task-specific metrics and Table C.2 shows the metrics for global structure preservation.
Table C.1: Macroscopic traffic prediction evaluation with LSTM and GRU encoders.
@@
Table C.2: Global structure preservation of LSTM and GRU encoders in macroscopic traffic prediction task.